Robert arrived at the LML site at 9:15. The collector had been unplugged so the doors didn’t open and no sample was collected. The top plate, drip plate, and the strand frames were removed for cleaning. The body was rinsed with DI water and the fog collector components were then cleaned with 409 followed by thorough rinsing with DI water. The system was plugged back in, keep an eye out to see if it happens again.